2/3x+12 - 4/3 = 2/x+4
Answer by stanbon(75887)

2/(3x+12) - 4/3 = 2/(x+4)
-------
Multiply thru by 3(x+4) to get:
2 -4(x+4) = 6
2 - 4x - 16 = 6
-4x-14=6
-4x=20
x=-5
--------
Cheers,
Stan H.